For each of the following equations connecting x and y, if the rate of change of x is 2units, find the rate of change of y at the given instant

a) y= 3/(2x-3)3, x=2

b) y= x3 + 2 , y=10


1
Expert's answer
2021-10-29T03:04:01-0400

given, the rate of change of x is 2units i.e.,"\frac{dx}{dt}=2" ."a)\\
y=\frac{3}{(2x-3)^3}, at\space x=2\\
\frac{dy}{dt}=3.(-3)(2x-3)^{-2}.(2\frac{dx}{dt})\\
\frac{dy}{dt}=(-9)(2\times2-3)^{-2}.(2\times2)\space(\text{Since, x=2 and rate of change of x is 2 units given.})\\
\frac{dy}{dt}=-36\\
\text{Therefore, the rate of change of y at the given instant is -36.}\\
b)\\
 y= x^3 + 2 , y=10 \\
\frac{dy}{dt}=3x^2.\frac{dx}{dt}\\
\frac{dy}{dt}=3(10)^2\times2\\
\frac{dy}{dt}=600\\
\text{Therefore, the rate of change of y at the given instant is 600.}\\"